solve by the elimination method
7r-8s=-21
8r+7s=89
a.the solution is?
b.there are infinitely many solutions
c.there is no solution

Answer by mananth(16949)

7r-8s =-21 .............1
8r+7s= 89 .............2

multiply (1)by 7
Multiply (2) by 8
49 r + -56 s = -147
64 r + 56 s = 712

Add the two equations

113 r = 565

/ 113

r = 5

plug value of r In (1)

7 r + -8 s = -21
35 + -8 s = -21
-8 s = -21 + -35
-8 s = -56
s = 7